# bibliography.bib

### v3

@@ Abstract.
An algebraic set $V\subset\mathbb{R}^{n}$ of dimension $d$ is a $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic set if it can be described, both globally and locally, by polynomial equations with rational coefficients. We prove that every nonsingular real algebraic set $V\subset\mathbb{R}^{n}$ with nonsingular algebraic subsets $\{V_{i}\}_{i=1}^{\ell}$ in general position is Nash diffeomorphic to a ${\mathbb{Q}}$-nonsingular $\mathbb{Q}$-algebraic set $V^{\prime}\subset\mathbb{R}^{m}$ with $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic subsets $\{V^{\prime}_{i}\}_{i=1}^{\ell}$ in general position and the Nash diffeomorphism $h:V\to V^{\prime}$ sends each $V_{i}$ to $V_{i}^{\prime}$, where $m:=\max\{n,2d+1\}$, if $V$ is compact, or $m:=n+2d+3$, otherwise. As a byproduct we obtain ${\mathbb{Q}}$-algebraic relative version of Nash-Tognoli theorem for compact manifolds. A key result in the proof is the description of ${\mathbb{Z}}/2{\mathbb{Z}}$-homological cycles of real Grassmannian manifolds by $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic representatives via an explicit desingularization of real embedded Schubert varieties.
@@ Introduction
Section 2 is devoted to the study of ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of real Grassmannians. It is well known that each real Grassmannian ${\mathbb{G}}_{m,n}$ of affine $m$-planes in ${\mathbb{R}}^{m+n}$ can be embedded as an algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$, see [BCR98, Theorem 3.4.4]. Let us identify ${\mathbb{G}}_{m,n}$ with the above algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$. It is also well known that incidence conditions with respect to a complete flag induce a finite cellular decomposition of ${\mathbb{G}}_{m,n}$ such that the Euclidean closure of each cell is an algebraic subset of ${\mathbb{G}}_{m,n}$, see [MS74]. The closures of these cells are called Schubert varieties associated to the complete flag and they generate the ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of each ${\mathbb{G}}_{m,n}$. The main result of this section is that, if we choose the complete flag $0\subset{\mathbb{R}}\subset{\mathbb{R}}^{2}\subset\dots\subset{\mathbb{R}}^{m+n}$ with respect to the standard coordinates of ${\mathbb{R}}^{m+n}$, each Schubert variety $\sigma_{\lambda}$ defined by incidence conditions, prescribed by a partition $\lambda$ of the $m\times n$ rectangle, with respect to the above complete flag admits a ${\mathbb{Q}}$-desingularization (see Definition 2.3). Latter result ensures that each ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of each real Grassmannian ${\mathbb{G}}_{m,n}$ has a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic representative. This fact is crucial to construct ${\mathbb{Q}}$-algebraic relative bordisms in Section 3. We point out that very general algorithms in resolution of singularities in characteristic 00 provide a resolution which is functorial and invariant under field extension, see as references [Wlo05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, mentioned algorithms are strongly non-constructive, whereas, in our explicit desingularization procedure for Schubert varieties, which is inspired by Zelevinski’s paper [Zel83], we conbinatorically control the equations defining the center of each blow-up of the resolution. We refer the interested reader to Remark 2.8 for more details on how to deduce the existence of a ${\mathbb{Q}}$-desingularization of ${\mathbb{Q}}$-algebraic sets from general results in resolution of singularities and why our procedure is preferable in the case of Schubert varieties.
@@ 1. Review on (real) ${\mathbb{Q}}$-algebraic geometry
In this subsection we briefly recall basics of real and complex ${\mathbb{Q}}$-algebraic geometry introduced in [FG, GSa].
@@
Moreover, if $L$ is algebraically closed, a deep application of Hilbert’s Nullstellensatz and Galois Theory in [FG] guarantees that ${\mathcal{I}}_{L}(V)={\mathcal{I}}_{K}(V)L[x]$, thus the ideal ${\mathcal{I}}_{K}(V)$ of $K[x]$ gives complete information about $V\subset L^{n}$. Latter equality holds as well when both $L$ and $K$ are real closed fields by the Tarski-Seidenberg principle. However, it is false for a general field extension $L|K$. As an example, if $L|K={\mathbb{R}}|{\mathbb{Q}}$ and $V:=\{x_{1}+\sqrt[3]{2}x_{2}=0\}=\{x_{1}^{3}+2x_{2}^{3}=0\}\subset{\mathbb{R}}^{2}$, then
${\mathcal{I}}_{\mathbb{R}}(V)=(x_{1}+\sqrt[3]{2}x_{2}){\mathbb{R}}[x_{1},x_{2}]\supsetneqq(x_{1}^{3}+2x_{2}^{3}){\mathbb{R}}[x_{1},x_{2}]={\mathcal{I}}_{\mathbb{Q}}(V){\mathbb{R}}[x_{1},x_{2}].$ (1.1)
@@
and by ${\mathcal{I}}_{\mathbb{Q}}(V)$ the vanishing ideal of $V$ in ${\mathbb{Q}}[x]$, as above. The following notion of ${\mathbb{R}}|{\mathbb{Q}}$-regular point was introduced in [FG].
@@
In [FG], the authors show that the set $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)$ is a non-empty Zariski open subset of $\operatorname{Reg}(V)$. It may happen that $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)\subsetneq\operatorname{Reg}(V)$. For instance, the ${\mathbb{Q}}$-algebraic line $V:=\{x_{1}+\sqrt[3]{2}x_{2}=0\}=\{x_{1}^{3}+2x_{2}^{3}=0\}\subset{\mathbb{R}}^{2}$ is regular (as an algebraic set), but $(0,0)$ is not ${\mathbb{R}}|{\mathbb{Q}}$-regular. This leads to the following definition originally introduced in [GSa, Definition 1.2].
@@
In [FG] the authors characterized the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity via a ${\mathbb{R}}|{\mathbb{Q}}$-jacobian criterion. In particular, if $V\subset{\mathbb{R}}^{n}$ of dimension $d$ is a nonsingular ${\mathbb{Q}}$-algebraic set, then $V$ is ${\mathbb{Q}}$-nonsingular if and only if for every $a\in V$ there are $q_{1},\dots,q_{n-d}\in\mathcal{I}_{\mathbb{Q}}(V)$ such that $\nabla q_{1}(a),\dots,\nabla q_{n-d}(a)$ are linearly independent and there exists a Zariski open subset $U$ of ${\mathbb{R}}^{n}$ such that $V\cap U=\mathcal{Z}(q_{1},\dots,q_{n-d})\cap U$. A systematic treatment on equivalent definitions of ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets involving Galois Theory, the complex Zariski closure of a (real) ${\mathbb{Q}}$-algebraic set and the vanishing ideal in the ring of global rational functions can be found in [Sav-T, §1.6].
Here we recall a crucial consequence of previous ${\mathbb{R}}|{\mathbb{Q}}$-jacobian criterion proved in [FG] and deeply applied in that paper. Its importance will be clear in the proof of our main theorem, more precisely in the proof of Lemma 4.1 below.
@@ 1.2. Fundamental examples
In [GSa, Lemmas 2.18 & 2.21], Ghiloni and the author proved that both ${\mathbb{G}}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}}$ and ${\mathbb{E}}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}+m+n}$ are projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets. Those algebraic sets are fundamental examples to represent cobordism classes of smooth manifolds and to develop algebraic approximation techniques over ${\mathbb{Q}}$ as in [GSa]. However, for the purposes of this paper we need further examples to develop relative techniques with equations over ${\mathbb{Q}}$ in the spirit of [AK81b].
@@
We prove that the polynomial $\mathrm{tr}(X)-m$ and the polynomial components of $\phi$ do suffice to describe nonsingular points of ${\mathbb{E}}^{*}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ via the ${\mathbb{Q}}$-jacobian criterion of [FG, Theorem 5.1.6]. As in the proof of [FG, Lemma 2.18], it suffices to show that $\mathrm{rnk}\,J_{\phi}(A,b,c)\geq(m+n)^{2}+m+n+1-m(n+1)-1=(m+n)^{2}-mn+n$ for all $(A,b,c)\in{\mathbb{E}}^{*}_{m,n}$.
@@
where $\mu(x)\in{\mathbb{G}}\subset{\mathbb{R}}^{(m+n)^{2}}$ is in matrix form. We prove that the polynomial components of $\phi$ do suffice to describe nonsingular points of $\mu^{*}({\mathbb{E}}^{*})\subset{\mathbb{R}}^{k}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ via the ${\mathbb{Q}}$-jacobian criterion of [FG]. As before, it suffices to show that $\mathrm{rnk}\,J_{\phi}(a,b,c)\geq k-d+n+1$ for all $(a,b,c)\in\mu^{*}({\mathbb{E}}^{*})$.
@@ 2.1. ${\mathbb{Q}}$-Desingularization of real embedded Schubert varieties
Let $V\subset{\mathbb{R}}^{m}$ be a ${\mathbb{Q}}$-algebraic set of dimension $d$. We say that $V^{\prime}\subset{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}$, for some $n\in{\mathbb{N}}$, is a desingularization of $V$ if $V^{\prime}\subset{\mathbb{R}}^{m+n}$ is a nonsingular algebraic set of dimension $d$ and $\pi|_{V^{\prime}}:V^{\prime}\to V$ is a birational map, where $\pi:{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is the projection onto the first factor. If, in addition, $V^{\prime}\subset{\mathbb{R}}^{m+n}$ is a ${\mathbb{Q}}$-nonsingular projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-algebraic we say that $V^{\prime}$ is a ${\mathbb{Q}}$-desingularization of $V$.
The goal of this section is to find ${\mathbb{Q}}$-desingularizations of embedded Schubert varieties defined by incidence conditions with respect to the canonical complete flag of ${\mathbb{R}}^{(m+n)^{2}}$, that is, to prove next result.
@@
In next result we adapt to our real embedded setting a desingularization technique developed by Zelevinsky in [Zel83].
@@
Then the algebraic set:
@@
By Remark 2.5, in order to prove Theorem 2.4 we are only left to prove next result.
@@
Each algebraic fibre bundle $Z_{\lambda}\subset{\mathbb{R}}^{(m+n)^{2}c}$ as in Lemma 2.6 is a projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set.
@@
By definition, $Z_{\lambda}$ is a ${\mathbb{Q}}$-algebraic subset of ${\mathbb{R}}^{(m+n)^{2}c}$ defined by the following equations in the variables $X:=(x_{i,j})_{i,j=1,\dots,m+n}$ and $Y_{k}:=(y^{(k)}_{i,j})_{i,j=1,\dots,m+n}$, for $k=1,\dots,c-1$:
@@
We point out that Theorem 2.4 follows as well by general algorithms for resolution of singularities, indeed the invariance under field extension is explicitly mentioned in [Wlo05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, there are some points to clarify. These algorithms in characteristic 00 actually work by resolving the marked ideal associated to an algebraic variety so, in our setting, we actually resolve the vanishing ideal $\mathcal{I}_{\mathbb{Q}}(V)$ of an affine ${\mathbb{Q}}$-algebraic set $V\subset{\mathbb{R}}^{n}$. We recall that in general $\mathcal{I}_{\mathbb{Q}}(V)\subset\mathcal{I}(V)$, so in principle the resolution will differ with respect to the classical Hironaka resolution for the algebraic set $V\subset{\mathbb{R}}^{n}$. The algorithm is claimed, in all mentioned references, to be stable under field extension. As explained in [Wlo05], a possible strategy is to deduce the invariance under field extension in three steps: first Włodarczyk deals with extensions of algebraically closed fields in Section 5.4, essentially by using Lefschetz principle, then the author proves the statement for the extension $\overline{K}|K$ of a field $K$ of characteristic 00, see Section 5.7, and finally in the same section Włodarczyk deduces the general result by a scheme theoretical argument. The invariance under algebraic closure in step two is not explicitly proved but it can be deduced by proving that, if a marked ideal $\underline{\mathcal{I}}$ over $\overline{K}$ is $K$-equivariant, then the derivative ideal $\mathcal{D}(\underline{\mathcal{I}})$ of $\underline{\mathcal{I}}$ over $\overline{K}$ is $K$-equivariant as well. Mentioned result is not obvious, indeed the derivative ideal of a marked ideal is locally defined so the local étale description of derivatives naturally involves coefficients over $\overline{K}$. However, the $K$-equivariance of $\mathcal{D}(\underline{\mathcal{I}})$, when $\underline{\mathcal{I}}$ is $K$-equivariant, can be proven by a computation applying Kähler differentials. The latter property allows us to run the algorithm locally preserving the coefficients of the involved equations over $K$. Then, the local centers at each stage of the resolution have to be globalized in such a way the center of each blow-up is globally a $K$-nonsingular $K$-algebraic set, obtaining finally the invariance under the field extension $\overline{K}|K$ by induction. The proof in [Kol07] uses a similar argument based on the theory of schemes and in [BM08] the result is only claimed. However, as previously mentioned, the application of scheme theoretical arguments is not really satisfying for a concrete point of view in the resolution procedure. We remark as well that our notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity turns out to be satisfied by the resulting real nonsingular algebraic set produced by the considered algorithms for resolution of singularities, see [Wlo05, Proposition 2.5.1] and [GSa, Corollary 2.4] for more details.
@@
As already mentioned, the algorithms proposed in [Wlo05, Kol07, BM08] are very non-constructive and the complexity of the resulting resolutions is very high since the procedure is local and it has to be globalized by comparing all the local resolutions. On the contrary, our algorithm based on Zelevinski’s resolution of singular Schubert varieties is much more elementary, very explicit and the complexity is controlled. Indeed, the number of steps to resolve the singularities of a Schubert variety turns out to be exactly the number of depressions of the associated Young diagram which is at most $c=\min\{m,n\}$ for each Schubert variety of ${\mathbb{G}}_{m,n}$. Hence, such a simple and explicit ${\mathbb{Q}}$-desingularization of real Schubert varieties has interest on its own.

### v4

# bibliography-2.bib

@@ Abstract.
An algebraic set $V\subset\mathbb{R}^{n}$ of dimension $d$ is a $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic set if it can be described, both globally and locally, by polynomial equations with rational coefficients. Set $m:=\max\{n,2d+1\}$, if $V$ is compact, or $m:=n+2d+3$, otherwise. We prove that every nonsingular real algebraic set $V\subset\mathbb{R}^{n}$ with nonsingular algebraic subsets $\{V_{i}\}_{i=1}^{\ell}$ in general position is Nash diffeomorphic to a $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic set $V^{\prime}\subset\mathbb{R}^{m}$ with $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic subsets $\{V^{\prime}_{i}\}_{i=1}^{\ell}$ in general position and the Nash diffeomorphism $h:V\to V^{\prime}$ sends each $V_{i}$ to $V_{i}^{\prime}$. As a byproduct we obtain a $\mathbb{Q}$-algebraic relative version of Nash-Tognoli theorem for compact manifolds. A key result in the proof is the description of $\mathbb{Z}/2\mathbb{Z}$-homological cycles of real Grassmannian manifolds by $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic representatives via the Bott-Samelson resolution of real embedded Schubert varieties.
@@ Introduction
Section 2 is devoted to the study of ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of real Grassmannians. It is well known that each real Grassmannian ${\mathbb{G}}_{m,n}$ of affine $m$-planes in ${\mathbb{R}}^{m+n}$ can be embedded as an algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$, see [BCR98, Theorem 3.4.4]. Let us identify ${\mathbb{G}}_{m,n}$ with the above algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$. It is also well known that incidence conditions with respect to a complete flag induce a finite cellular decomposition of ${\mathbb{G}}_{m,n}$ such that the Euclidean closure of each cell is an algebraic subset of ${\mathbb{G}}_{m,n}$, see [MS74]. The closures of these cells are called Schubert varieties associated to the complete flag and they generate the ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of each ${\mathbb{G}}_{m,n}$. The main result of this section is that, if we choose the complete flag $0\subset{\mathbb{R}}\subset{\mathbb{R}}^{2}\subset\dots\subset{\mathbb{R}}^{m+n}$ with respect to the standard coordinates of ${\mathbb{R}}^{m+n}$, each Schubert variety $\sigma_{\lambda}$ defined by incidence conditions, prescribed by a partition $\lambda$, with respect to the above complete flag admits a ${\mathbb{Q}}$-desingularization (see Definition 2.3). Latter result ensures that each ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of each real Grassmannian ${\mathbb{G}}_{m,n}$ has a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic representative. This is a key property for the construction of ${\mathbb{Q}}$-algebraic relative bordisms in Section 3. We point out that very general algorithms in resolution of singularities in characteristic 00 provide a resolution which is functorial and invariant under field extension, see as references [Wlo05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, mentioned algorithms have a very high complexity, thus it is a very hard task in general to concretely apply them obtaining an explicit resolution of a given algebraic variety, see [FKP]. By contrast, the application of the very explicit Bott-Samelson desingularization procedure for real embedded Schubert varieties, see [Dem74, Zel83] and the books [FP98, Man01, Bri05], allows us to conbinatorically control the equations defining the center of each blow-up of the resolution. We emphasize that, to the best of the author’s knowledge, the ${\mathbb{R}}|{\mathbb{Q}}$-regularity of real embedded Bott-Samelson varieties is not a direct consequence of results already known in the literature, although Demazure [Dem74] develops the desingularization procedure in the very general setting of flag varieties over a field $k$. Indeed, for our purposes, the choice of an appropriate embedding in some real affine space of each Grassmannian manifold ${\mathbb{G}}_{m,n}$ and of the resulting desingularization of its embedded Schubert subvarieties plays a crucial role for the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity, see Example 1.3 for a clear explanation. We refer the interested reader to Remark 2.8 for more details on how to deduce the existence of a ${\mathbb{Q}}$-desingularization of ${\mathbb{Q}}$-algebraic sets from general results in resolution of singularities and why Bott-Samelson procedure is preferable in the case of Schubert varieties.
@@ 1. Review on (real) ${\mathbb{Q}}$-algebraic geometry
In this subsection we briefly recall some fundamental notions of real and complex ${\mathbb{Q}}$-algebraic geometry introduced in [FG, GSa].
@@
Moreover, if $L$ is algebraically closed, an application of Hilbert’s Nullstellensatz and Galois Theory in [FG, Theorem 1.3.4] guarantees that ${\mathcal{I}}_{L}(V)={\mathcal{I}}_{K}(V)L[x]$, thus the ideal ${\mathcal{I}}_{K}(V)$ of $K[x]$ gives complete information about $V\subset L^{n}$. Latter equality holds as well when both $L$ and $K$ are real closed fields by the Tarski-Seidenberg principle, see [FG, Lemma 1.4.1]. However, this is false for a general field extension $L|K$. As an example, if $L|K={\mathbb{R}}|{\mathbb{Q}}$ and $V:=\{x-\sqrt[3]{2}=0\}=\{x^{3}-2=0\}\subset{\mathbb{R}}$, then
${\mathcal{I}}_{\mathbb{R}}(V)=(x-\sqrt[3]{2}){\mathbb{R}}[x]\supsetneqq(x^{3}-2){\mathbb{R}}[x]={\mathcal{I}}_{\mathbb{Q}}(V){\mathbb{R}}[x].$ (1.1)
@@
and by ${\mathcal{I}}_{\mathbb{Q}}(V)$ the vanishing ideal of $V$ in ${\mathbb{Q}}[x]$, as above. The following notion of ${\mathbb{R}}|{\mathbb{Q}}$-regular point was introduced in [FG, Definition 4.1.1].
@@
In [FG, Theorem 4.4.1], Fernando and Ghiloni show that the set $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)$ is a non-empty Zariski open subset of $\operatorname{Reg}(V)$. Though, it may happen that $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)\subsetneq\operatorname{Reg}(V)$, see the example below.
Example 1.3.
[…]
This leads to the following definition originally introduced in [GSa, Definition 1.2].
@@
In [FG, Theorem 4.1.7] the authors characterized the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity via a ${\mathbb{R}}|{\mathbb{Q}}$-Jacobian criterion. In particular, if $V\subset{\mathbb{R}}^{n}$ of dimension $d$ is a nonsingular ${\mathbb{Q}}$-algebraic set, then $V$ is ${\mathbb{Q}}$-nonsingular if and only if for every $a\in V$ there are $q_{1},\dots,q_{n-d}\in\mathcal{I}_{\mathbb{Q}}(V)$ such that $\nabla q_{1}(a),\dots,\nabla q_{n-d}(a)$ are linearly independent and there exists a Zariski open subset $U$ of ${\mathbb{R}}^{n}$ such that $V\cap U=\mathcal{Z}(q_{1},\dots,q_{n-d})\cap U$. A systematic treatment on equivalent definitions of ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets involving Galois Theory, the complex Zariski closure of a (real) ${\mathbb{Q}}$-algebraic set and the vanishing ideal in the ring of global rational functions can be found in [Sav-T, §1.6].
Here we recall a crucial consequence of mentioned ${\mathbb{R}}|{\mathbb{Q}}$-Jacobian criterion of [FG] and deeply applied in that paper. Its importance will be clear in the proof of our Main Theorem, more precisely, in the proof of Lemma 4.1 below.
@@ 1.2. Fundamental examples
In [GSa, Lemmas 2.18 & 2.21], Ghiloni and the author proved that both ${\mathbb{G}}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}}$ and ${\mathbb{E}}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}+m+n}$ are projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets. These algebraic sets are fundamental examples to represent cobordism classes of smooth manifolds and to develop algebraic approximation techniques over ${\mathbb{Q}}$ as in [GSa]. However, for the purposes of this paper we need further examples to develop relative techniques ‘over ${\mathbb{Q}}$’ in the spirit of [AK81b].
@@
We prove that the polynomial $\mathrm{tr}(X)-m$ and the polynomial components of $\phi$ do suffice to describe nonsingular points of ${\mathbb{E}}^{*}_{m,n}\subset{\mathbb{R}}^{(m+n)^{2}}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ via the ${\mathbb{R}}|{\mathbb{Q}}$-Jacobian criterion [FG, Theorem 4.1.7]. As in the proof of [GSa, Lemma 2.18], it suffices to show that $\mathrm{rnk}\,J_{\phi}(A,b,c)\geq(m+n)^{2}+m+n+1-m(n+1)-1=(m+n)^{2}-mn+n$ for all $(A,b,c)\in{\mathbb{E}}^{*}_{m,n}$.
@@
where $\mu(x)\in{\mathbb{G}}\subset{\mathbb{R}}^{(m+n)^{2}}$ is in matrix form. We prove that the polynomial components of $\phi$ do suffice to describe nonsingular points of $\mu^{*}({\mathbb{E}}^{*})\subset{\mathbb{R}}^{k}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ via the ${\mathbb{R}}|{\mathbb{Q}}$-Jacobian criterion [FG, Theorem 4.1.7]. As before, it suffices to show that $\mathrm{rnk}\,J_{\phi}(a,b,c)\geq k-d+n+1$ for all $(a,b,c)\in\mu^{*}({\mathbb{E}}^{*})$.
@@ 2.1. ${\mathbb{Q}}$-Desingularization of real embedded Schubert varieties
Let $V\subset{\mathbb{R}}^{m}$ be a ${\mathbb{Q}}$-algebraic set of dimension $d$. We say that $V^{\prime}\subset{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}$, for some $n\in{\mathbb{N}}$, is a desingularization of $V$ if $V^{\prime}\subset{\mathbb{R}}^{m+n}$ is a nonsingular algebraic set of dimension $d$ and $\pi|_{V^{\prime}}:V^{\prime}\to V$ is a birational map, where $\pi:{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is the projection onto the first factor. If, in addition, $V^{\prime}\subset{\mathbb{R}}^{m+n}$ is a ${\mathbb{Q}}$-nonsingular projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-algebraic set we say that $V^{\prime}$ is a ${\mathbb{Q}}$-desingularization of $V$.
It is very well-known since Demazure [Dem74] that Bott-Samelson method produces a desingularization of Schubert varieties. This algorithm has been deeply studied in literature, for instance Zelevinsky [Zel83] proved that a precise choice on the order of the subvarieties to blow-up produces a small resolution. The goal of this section is to show that the Bott-Samelson method produces ${\mathbb{Q}}$-desingularizations of embedded Schubert varieties defined by incidence conditions with respect to the canonical complete flag of ${\mathbb{R}}^{m+n}$, that is, to prove the next result.
@@
In next result we translate to our real embedded setting the Bott-Samelson desingularization technique in [Dem74, Zel83].
@@
Then the embedded Bott-Samelson algebraic set:
@@
By Remark 2.5, in order to prove Theorem 2.4 we are only left to prove that the explicit equations of each embedded Bott-Samelson algebraic set $Z_{\lambda}$ of Lemma 2.6 actually describe the local smooth behavior of $Z_{\lambda}$ at any point as well.
@@
Each embedded Bott-Samelson algebraic set $Z_{\lambda}\subset{\mathbb{R}}^{(m+n)^{2}c}$ as in Lemma 2.6 is a projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set.
@@
By definition, $Z_{\lambda}$ is a ${\mathbb{Q}}$-algebraic subset of ${\mathbb{R}}^{(m+n)^{2}c}$ defined by the following equations in the variables $X:=(x_{ij})_{i,j=1,\dots,m+n}$ and $Y_{k}:=(y^{(k)}_{ij})_{i,j=1,\dots,m+n}$, for $k=1,\dots,c-1$:
@@
We point out that Theorem 2.4 follows as well by general algorithms for resolution of singularities, indeed the invariance under field extension is explicitly mentioned in [Wlo05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, there are some points to clarify. These algorithms in characteristic 00 actually work by resolving the marked ideal associated to an algebraic variety so, in our setting, we actually resolve the vanishing ideal $\mathcal{I}_{\mathbb{Q}}(V)$ of an affine ${\mathbb{Q}}$-algebraic set $V\subset{\mathbb{R}}^{n}$. We recall that in general $\mathcal{I}_{\mathbb{Q}}(V)\subset\mathcal{I}(V)$, so in principle the resolution will differ with respect to the classical Hironaka resolution for the algebraic set $V\subset{\mathbb{R}}^{n}$. The algorithm is claimed, in all mentioned references, to be stable under field extension. As explained in [Wlo05], a possible strategy is to deduce the invariance under field extension in three steps: first Włodarczyk deals with extensions of algebraically closed fields in Section 5.4, essentially by using Chevalley’s Theorem as a transfer principle, then the author proves the statement for the extension $\overline{K}|K$ of a field $K$ of characteristic 00, see Section 5.7, and finally in the same section Włodarczyk deduces the general result by a scheme theoretical argument. The invariance under algebraic closure in step two is not explicitly proved but it can be deduced by proving that, if a marked ideal $\underline{\mathcal{I}}$ over $\overline{K}$ is $K$-equivariant, then the derivative ideal $\mathcal{D}(\underline{\mathcal{I}})$ of $\underline{\mathcal{I}}$ over $\overline{K}$ is $K$-equivariant as well. Mentioned result is not obvious, indeed the derivative ideal of a marked ideal is locally defined so the local étale description of derivatives naturally involves coefficients over $\overline{K}$. However, the $K$-equivariance of $\mathcal{D}(\underline{\mathcal{I}})$, when $\underline{\mathcal{I}}$ is $K$-equivariant, can be proven by a computation applying Kähler differentials. The latter property allows us to run the algorithm locally preserving the coefficients of the involved equations over $K$. Then, the local centers at each stage of the resolution have to be globalized in such a way the center of each blow-up is globally a $K$-nonsingular $K$-algebraic set, obtaining finally the invariance under the field extension $\overline{K}|K$ by induction. The proof in [Kol07] uses a similar argument based on the theory of schemes and in [BM08] the result is only claimed. However, as previously mentioned, the application of scheme theoretical arguments is not really satisfying for a concrete point of view in the resolution procedure. We remark as well that our notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity turns out to be satisfied by the resulting real nonsingular algebraic set produced by the considered algorithms for resolution of singularities, see [Wlo05, Proposition 2.5.1] and [GSa, Corollary 2.4] for more details.
@@
As already mentioned, the complexity of the algorithms proposed in [Wlo05, Kol07, BM08] is very high since the procedure is local and it has to be globalized by comparing all the local resolutions so it becomes very hard in practice to apply the algorithm to a given algebraic set, see [FKP] for more details. On the contrary, Bott-Samelson resolution of singular Schubert varieties is much more elementary, very explicit and the complexity is controlled. Indeed, the number of steps to resolve the singularities of a Schubert variety turns out to be exactly the number of depressions of the associated Young diagram, which is at most $c=\min\{m,n\}$ for each Schubert variety of ${\mathbb{G}}_{m,n}$. Hence, such a simple and explicit ${\mathbb{Q}}$-desingularization of real Schubert varieties has interest on its own.